# Tracing Nitrogen Enrichment across Cosmic Time with JWST

### v2

@@ 2.1.2 JWST Archival Sample
We compiled a sample of 11 galaxies at z>z>1 observed with JWST from the literature, all showing robust detections (S/N ¿ 3) of the [O iii] $\lambda$4363 auroral line, of [O ii] $\lambda\lambda$3727,3729 and of [N ii] $\lambda$6585 (Sanders et al., 2024; Welch et al., 2024; Arellano-Córdova et al., 2025b; Welch et al., 2025; Zhang et al., 2025). Two of these objects also exhibits detections of the [O ii] $\lambda\lambda$7320,7330 auroral doublet. An additional 55 galaxies meeting the same selection criteria were drawn from version 4 of the DAWN JWST Archive (DJA; Valentino et al., 2025), a public repository containing reduced spectroscopic data from archival JWST observations processed with MSAExp (Brammer, 2023). The sample selection was based on the detection of optical [N ii] lines, while nitrogen UV lines were not considered, to avoid potential systematic effects due to the different tracers.
@@ 2.1.3 Local Universe sample
For comparison with the sample presented in the previous section, we assembled a large sample of low-redshift galaxies and local H ii regions, with robust auroral-line detections, enabling direct metallicity determinations, for a total of 904 local objects. All abundances were re-derived in a self-consistent manner, by recovering the published line fluxes and re-computing electron temperatures following a uniform methodology. A conservative cut of ${\rm S/N}>3$ on the relevant auroral lines was applied throughout. The local compilation includes:
@@ 2.2.1 $z>1$ sample
We included galaxies from the ground-based MOSDEF (MOSFIRE Deep Evolution Field survey, Kriek et al., 2015) and KLEVER (KMOS Lensed Emission-Line Velocity and Metallicity Survey, Curti et al., 2020) surveys. These surveys do not detect auroral lines but provide all the strong emission line fluxes required to derive metallicities and nitrogen abundances through indirect methods, in particular [O ii] $\lambda\lambda$3727,3729 and [N ii] $\lambda$6585. MOSDEF (PI: M. Kiek) is a spectroscopic survey carried out with MOSFIRE (McLean et al., 2012) on the 10m Keck I telescope, targeting galaxies at $1.4\leq z\leq 3.8$. KLEVER, on the other hand, is an ESO Large Programme (PID: 197.A-0717, PI: M. Cirasuolo) that observed 192 galaxies in the redshift range $1.2<z<2.5$ with KMOS. We applied a uniform and conservative cut requiring ${\rm S/N}\geq 5$ on all emission lines of interest.
We also considered galaxies from the public release of the JWST Advanced Deep Extragalactic Survey (JADES) Data Release 4 (Robertson et al., 2023; Eisenstein et al., 2023; Curtis-Lake et al., 2025; Scholtz et al., 2025). From this catalogue, we selected sources at $z>1$ without auroral-line detections (as those with detected auroral lines are already included in the high-redshift sample described above), but with robust measurements of the strong nebular lines relevant to our analysis. We applied the same S/N threshold (${\rm S/N}\geq 3$ on key strong lines) and adopted standard diagnostics from Kewley et al. (2013) to exclude active galactic nuclei (AGN) contamination. This selection yielded a final sample of 157 galaxies from JADES DR4 suitable for inclusion in our analysis. Finally, we complemented this dataset with additional galaxies drawn from version 4 of the DAWN JWST Archive selected following the same criteria, as described in Section 2.1.2, but without the auroral-line requirement. This brings the total number of high-redshift, strong-line–based comparison galaxies to 430. The distribution of our high-redshift galaxies is consistent with the star-forming locus, with no clear evidence for contamination by AGN or LINERs (low-ionisation nuclear emission-Line region).
@@ 2.2.2 Local Universe sample
To provide a comparison with a sample representative of the local galaxy population and extend the local comparison sample to high metallicity we leverage the SDSS Data Release 8 MPA-JHU catalogue (Brinchmann et al., 2004). We applied selection cuts requiring a S/N ratio greater than five on key emission lines, restricted the sample to local redshifts ($z<0.27$), and excluded AGN contamination based on the BPT classification criteria of Kewley et al. (2013). The final SDSS sample includes approximately 87,000 galaxies.
@@ 3.3 Oxygen abundances from strong lines
Figure 3: Relation between the N/O ratio measured using the $T_{e}$ method and the N2O2 (left) and N2S2 (right) diagnostics for our full high-$z$ auroral-line sample. For comparison, we include local galaxies with auroral-line detections from the literature as gray points (see references in Sec. 2.1.3). The local calibration from Pérez-Montero & Contini (2009) is shown in blue and a fit to the SDSS stacks of Curti et al. (2017) in green. In both panels, the grey error bar in the top-left corner represents the median measurement uncertainty of the local comparison sample. The red solid line shows the best fit obtained from the $z\sim$2-4 sample with the uncertainty range shown as a shading.
@@ 3.4 Nitrogen abundances from direct measurements
For the galaxies with a direct electron temperature measurement, nitrogen abundances were computed from the dust-corrected fluxes of the [N ii] $\lambda 6584$ and [O ii] $\lambda\lambda 3727,3729$ emission lines. We assumed that the ionic abundances of N+ and O+ trace the nitrogen and oxygen content in the low-ionisation zone of the nebula, and adopted the same electron temperature ($T_{2}$) for both ions. This choice is justified by the similar ionisation potentials of N+ (29.6 eV) and O+ (35.1 eV), which ensures that they originate from roughly the same physical region (e.g., Garnett, 1990; Vila Costas & Edmunds, 1993).
As detailed in Sec. 3.2 we measure $T_{2}$ directly when possible. Alternatively we infer it from the $T_{2}$$T_{3}$ relation of C25. We also tested how the results would change when using $T_{e}$([N ii]) estimated from the empirical relation between $T_{e}$([N ii]) and $T_{e}$([O iii]) available in the literature. This relation is well-calibrated and relatively tight in the low-redshift Universe (no auroral lines from [N ii] are available for the high-$z$ sample), as shown in previous studies (e.g., Garnett, 1992; Pérez-Montero & Contini, 2009). When adopting $T_{e}$([N ii]) for N+ and $T_{e}$([O ii]) for O+ the derived log(N/O) increases by $\sim$ 0.06 dex.
@@
While it is true that such ICF prescriptions could evolve with cosmic time, we have tested whether their application introduces artificial trends in our results by repeating all analyses without applying any ionisation correction. The ICF values in our sample are typically modest, up to 1.2 with a median value of $\sim$1.1. We find that removing the ICF corrections does not alter any of the conclusions presented in this work, introducing at most a small systematic offset (0.07 dex higher) in the absolute N/O values. This demonstrates that our results are not affected by artifacts introduced by potential ICF redshift evolution.
@@ 4.1 Strong-line diagnostics for N/O at high redshift
As shown in Figure 3, the high-redshift galaxies in our sample occupy the same locus as local sources and broadly follow the trend defined by previous empirical calibrations, with no large systematic offset. However, they tend to lie toward the upper envelope of the local distribution, with slightly elevated N/O values with respect to the local distribution at fixed N2O2 or N2S2. To quantify their behaviour, we performed a linear regression for both the diagnostics. The regressions were performed on the zz¡4 subsample of objects to ensure internal consistency and focus on systems around Cosmic Noon, but we verified that expanding it to the whole zz¿1 sample gives minimal changes in the computed coefficients, in both cases within their uncertainties. The best-fit relations were obtained using a Huber regressor, a robust algorithm that combines the properties of ordinary least-squares and absolute-deviation fitting. It behaves like least-squares for well-behaved data points but progressively downweights outliers beyond a certain threshold, providing stable estimates even in the presence of intrinsic scatter or uncertain measurements-conditions typical of high-redshift spectroscopic observations. Uncertainties on the slope and intercept were estimated via the jackknife resampling technique, which recomputes the fit after iteratively removing one data point at a time. This method is preferred over bootstrap resampling given the relatively small size of the high-z sample, which would otherwise lead to excessive repetition of individual objects across subsamples.
Our fiducial best-fit relation on the zz¡4 objects is then
$\log(\mathrm{N/O})=(0.87\pm 0.10)\times\mathrm{N2O2}-(0.20\pm 0.11).$ (3)
The intrinsic scatter of this relation, estimated after accounting for measurement uncertainties, is $\sigma_{\mathrm{int}}=0.09$ dex. The resulting best-fit relation is therefore slightly steeper than that derived from SDSS stacks, reflecting the tendency of our high-redshift galaxies to lie along the upper envelope of the local distribution. Nevertheless, the coefficients remains consistent within the uncertainties with the local calibration of Pérez-Montero & Contini (2009).
A similar conclusion holds for the N2S2–N/O relation (right panel of Fig. 3), despite its slightly larger intrinsic scatter of 0.13 dex compared to N2O2: the high-redshift galaxies – both from the MARTA sample and from the literature – are consistent with the distribution of local galaxies and the calibrations derived in the nearby Universe. The newly derived empirical fit are in good agreement with local trends, supporting the use of N2S2 as a secondary tracer of N/O in distant galaxies. For the zz¡4 dataset, the best-fit N2S2–N/O relation is
$\log(\mathrm{N/O})=(0.82\pm 0.19)\times\mathrm{N2S2}-(0.94\pm 0.05).$ (4)
@@
Figure 4: Relation between N/O and O/H for our full sample, as described in Section 2. Gold hexagons indicate MARTA galaxies with direct measurements, and light-blue triangles show the JWST auroral-line dataset. High-redshift strong-line galaxies from the literature are plotted as teal triangles, and local galaxies (both direct and strong-line measurements) as grey circles. White squares represent local galaxies binned in metallicity, showing the median N/O value in each bin. Two violet stars indicate the median N/O values for high-redshift galaxies, with the first corresponding to $12+\log(\mathrm{O/H})<8.1$ and the second to $12+\log(\mathrm{O/H})\geq 8.1$. The black line shows the best-fit relation derived from the binned values of local galaxies using the functional form from Hayden-Pawson et al. (2022), with the best-fit parameters obtained in this work and reported in Table 1. The solar abundance values from Asplund et al. (2021) are shown with the orange Sun symbol. Finally, the inset panel presents the histogram of N/O residuals relative to the low-redshift fit, separately for local and high-redshift galaxies, with vertical dashed lines marking the medians of the two distributions.
@@ 4.2 N/O-O/H relationship
Table 1: The best-fit parameters of the median N/O–O/H relation for the local galaxy sample are listed, using equation 5. Here, $12+\log(\mathrm{O/H})_{0}=8.60$ corresponds to $(\mathrm{O/H})_{0}=4\times 10^{-4}$.
@@
$1.385\pm 0.017$ $8.60\pm 0.21$ $2.8\pm 0.6$ $2.3\pm 0.8$
Figure 4 demonstrates a systematic nitrogen enhancement of the high-$z$ galaxy population relative to the local relation, particularly at lower metallicities. This is highlighted by the median N/O points as measured in the two different O/H regimes, since both lie above the local fit. The inset panel in the upper left corner shows the distribution of residuals relative to the local best-fit relation and further demonstrates that the high-$z$ sample (teal histogram) is shifted toward positive values compared to the local sample, in grey.
To quantitatively assess the difference between the two distributions, we compared the residuals of the high-$z$ and local samples (see Appendix C). The two distributions are statistically inconsistent with being drawn from the same parent population: the Anderson-Darling test [^2] The Anderson–Darling test is a non parametric test that, similarly to the Kolmogorov–Smirnov, compares the cumulative distributions of two samples. Unlike K-S, it assigns more weight to differences in the tails of the distributions and provides a more uniform sensitivity across quantiles. This makes it particularly suitable when the samples have very different sizes or when discrepancies in the extreme values are scientifically relevant. (AD-test, Scholz & Stephens, 1987) rejects the null hypothesis at high significance ($p<10^{-3}$). Additional tests and a full discussion are presented in Appendix C.
@@
We mostly confirm conclusions from previous works regarding nitrogen enrichment. Systems known to exhibit elevated N/O ratios (e.g. the Sunburst arc, the Zhang et al. (2025) source, EXCELS objects) lie above the local trend, and in some cases even lie outside the dispersion of our high-redshift trend. Galaxies reported to follow the local pattern (CECILIA D40, SGAS 1732) fall within the dispersion of the local population. Stiavelli et al. (2025) classify only three of the CEERS sources as N-enhanced, noting that the remaining two are consistent with the dispersion of the local relation. With our larger statistics we find all CEERS sources to be N-enhanced with respect to the local sample, and two of them to lie above the dispersion of the high-redshift one. Our findings therefore provide context for previous discussions of individual sources in the literature, confirming the existence of both N-enhanced and non-enhanced sources, and showing how their frequency changes with metallicity.
Figure 6: Same as Figure 4, but with the SDSS density contours as the sole local reference for simplicity. The high-$z$ sample is represented with circles, with no distinction between different datasets. Left panel: the high-$z$ sample is colour-coded by $\Delta\text{MS}$, i.e. the deviation from the parametrised SFMS from Popesso et al. (2023), assuming a reference redshift of $z=3$. Most of N-enhanced galaxies lie above the MS. Right panel: the high-$z$ sample is colour-coded by stellar mass. The gradient angle of the $M_{*}$ trend within the plane is shown as a red arrow, visually representing the finding that N-enhanced galaxies are preferentially associated with lower stellar masses.

### v3

@@ 2.1.2 JWST archival sample
We compiled a sample of 11 galaxies at z>z>1 observed with JWST from the literature meeting all the selection criteria required for auroral-lines sample. Two of these objects also exhibits detections of the [O ii] $\lambda\lambda$7320,7330 auroral doublet. An additional 66 galaxies meeting the same selection criteria were drawn from version 4 of the DAWN JWST Archive (DJA; Valentino et al., 2025), a public repository containing reduced spectroscopic data from archival JWST observations processed with MSAExp (Brammer, 2023). Emission line fluxes for these galaxies were adopted directly from the standard DJA catalogue. The sample selection was based on the detection of optical [N ii] lines, while nitrogen UV lines were not considered, to avoid potential systematic effects due to the different tracers.
@@ 2.1.3 Local universe sample
For comparison with the sample presented in the previous section, we assembled a large sample of low-redshift galaxies and local H ii regions, meeting all the S/N requirements of the auroral-line sample, enabling direct metallicity determinations, for a total of 904 local objects. All abundances were re-derived in a self-consistent manner, by recovering the published line fluxes and re-computing electron temperatures following a uniform methodology. The local compilation includes:
@@ 2.2.1 $z>1$ sample
We included galaxies from the ground-based MOSDEF (MOSFIRE Deep Evolution Field survey, Kriek et al., 2015) and KLEVER (KMOS Lensed Emission-Line Velocity and Metallicity Survey, Curti et al., 2020) surveys. These surveys do not detect auroral lines but provide all the strong emission line fluxes required to derive metallicities and nitrogen abundances through indirect methods, in particular [O ii] $\lambda\lambda$3727,3729 and [N ii] $\lambda$6585. MOSDEF (PI: M. Kiek) is a spectroscopic survey carried out with MOSFIRE (McLean et al., 2012) on the 10m Keck I telescope, targeting galaxies at $1.4\leq z\leq 3.8$. KLEVER, on the other hand, is an ESO Large Programme (PID: 197.A-0717, PI: M. Cirasuolo) that observed 192 galaxies in the redshift range $1.2<z<2.5$ with KMOS. We applied the S/N selection criteria described at the beginning of this section.
We also considered galaxies from the public release of the JWST Advanced Deep Extragalactic Survey (JADES) Data Release 4 (Robertson et al., 2023; Eisenstein et al., 2023; Curtis-Lake et al., 2025; Scholtz et al., 2025). From this catalogue, we selected sources at $z>1$ without auroral-line detections (as those with detected auroral lines are already included in the high-redshift sample described above), but with robust measurements of the strong nebular lines relevant to our analysis. This selection yielded a final sample of 188 galaxies from JADES DR4 suitable for inclusion in our analysis. Finally, we complemented this dataset with additional galaxies drawn from version 4 of the DAWN JWST Archive selected following the same criteria, as described in Section 2.1.2, but without the auroral-line requirement. This brings the total number of high-redshift, strong-line–based comparison galaxies to 535. The distribution of our high-redshift galaxies is consistent with the star-forming locus, with no clear evidence for contamination by AGN or LINERs (low-ionisation nuclear emission-Line region).
@@ 2.2.2 Local universe sample
To provide a comparison with a sample representative of the local galaxy population and extend the local comparison sample to high metallicity we leverage the SDSS Data Release 8 MPA-JHU catalogue (Brinchmann et al., 2004). In addition to the S/N selection criteria applied uniformly across all datasets, we restricted this sample to local redshifts ($z<0.27$). The final SDSS sample includes approximately 87,000 galaxies.
@@ 3.1 Dust reddening correction
We also tested the impact of adopting a high-redshift nebular attenuation curve as recently proposed by Reddy et al. (2026), using their ‘median’ attenuation law (Eq. 22). The two laws yield systematically different E(BV)E(B-V) values, with the Reddy et al. (2026) curve returning generally lower estimates compared to Cardelli et al. (1989), with a median offset of 0.06-0.06 dex. This translates into typical variations of $\sim+0.03$ dex in N2O2, a key quantity relevant for our conclusions, and of similar order in other properties (e.g., $\sim 50$ K in electron temperatures). Importantly, these differences do not affect our overall conclusions.
@@ 3.3 Oxygen abundances from strong lines
Figure 3: Relation between the N/O ratio measured using the $T_{e}$ method and the N2O2 (upper panel) and N2S2 (lower panel) diagnostics for our full high-$z$ auroral-line sample. For comparison, we include local galaxies with auroral-line detections from the literature as gray points (see references in Sec. 2.1.3). The local calibration from Pérez-Montero & Contini (2009) is shown in blue and a fit to the SDSS stacks of Curti et al. (2017) in green. In both panels, the grey error bar in the top-left corner represents the median measurement uncertainty of the local comparison sample. The red solid line shows the best fit obtained from the high-z sample with the uncertainty range shown as a shading.
@@ 3.4 Nitrogen abundances from direct measurements
For the galaxies with a direct electron temperature measurement, nitrogen abundances were computed from the dust-corrected fluxes of the [N ii] $\lambda 6584$ and [O ii] $\lambda\lambda 3727,3729$ emission lines. We assumed that the ionic abundances of N+ and O+ trace the nitrogen and oxygen content in the low-ionisation zone of the nebula, and adopted the same electron temperature ($T_{2}$) for both ions. This choice is justified by the similar ionisation potentials required both to form these ions from their neutral states (14.5 eV for $^{0}\rightarrow N$+ and 13.6 eV for $^{0}\rightarrow O$+) and to ionise them further (29.6 eV for $^{+}\rightarrow N$2+ and 35.1 eV for $^{+}\rightarrow O$2+), ensuring they originate from roughly the same physical region (e.g., Garnett, 1990; Vila Costas & Edmunds, 1993).
As detailed in Sec. 3.2 we measure $T_{2}$ directly when possible. This was feasible for only a small subset of the high-redshift sample (9 objects), and for $\sim$900 local Universe galaxies. Alternatively we infer it from the $T_{2}$$T_{3}$ relation of C25. We also tested how the results would change when using $T_{e}$([N ii]) estimated from the empirical relation between $T_{e}$([N ii]) and $T_{e}$([O iii]) available in the literature. This relation is well-calibrated and relatively tight in the low-redshift Universe (no auroral lines from [N ii] are available for the high-$z$ sample), as shown in previous studies (e.g., Garnett, 1992; Pérez-Montero & Contini, 2009). When adopting $T_{e}$([N ii]) for N+ and $T_{e}$([O ii]) for O+ the derived log(N/O) increases by $\sim$ 0.06 dex.
@@
While it is true that such ICF prescriptions could evolve with cosmic time, we have tested whether their application introduces artificial trends in our results by repeating all analyses without applying any ionisation correction. This latter approach is equivalent to assuming N+/O+ = N/O, which is the ICF prescription proposed by Peimbert & Costero (1969) based on the similar ionisation energies of N and O, and is commonly adopted in the literature when deriving N/O from singly ionised species. The ICF values in our sample are typically modest, up to 1.2 with a median value of $\sim$1.1. We find that removing the ICF corrections does not alter any of the conclusions presented in this work, introducing at most a small systematic offset (0.07 dex higher) in the absolute N/O values. This demonstrates that our results are not affected by artifacts introduced by potential ICF redshift evolution.
@@ 4.1 Strong-line diagnostics for N/O at high redshift
As shown in Figure 3, the high-redshift galaxies in our sample lie within the locus defined by local sources and are broadly consistent with the trend from previous empirical calibrations. However, while they formally fall within the scatter of the local distribution, in the N2O2 diagram, high-z objects predominantly occupy the upper envelope of the distribution, showing systematically elevated N/O values at fixed N2O2 and hinting at a possible redshift evolution of this relation. Conversely, in the N2S2 diagram they appear more scattered throughout the local distribution and exhibit better overall consistency with nearby galaxies.
To quantify their behaviour, we performed a linear regression for both the diagnostics. The regressions were performed on the whole $z>1$ sample. For consistency, we verified that splitting the sample in two redshift ranges ($1<z<4$; $4<z<6$) gives comparable coefficients within their uncertainties (see also Appendix B for further analysis). The best-fit relations were obtained using a Huber regressor, a robust algorithm that combines the properties of ordinary least-squares and absolute-deviation fitting. It behaves like least-squares for well-behaved data points but progressively downweights outliers beyond a certain threshold, providing stable estimates even in the presence of intrinsic scatter or uncertain measurements-conditions typical of high-redshift spectroscopic observations. Uncertainties on the slope and intercept were estimated via the jackknife resampling technique, which recomputes the fit after iteratively removing one data point at a time. This method is preferred over bootstrap resampling given the relatively small size of the high-z sample, which would otherwise lead to excessive repetition of individual objects across subsamples.
Our fiducial best-fit relation is then
$\log(\mathrm{N/O})=(1.02\pm 0.09)\times\mathrm{N2O2}-(0.03\pm 0.06).$ (3)
The estimated intrinsic scatter of this relation is $\sigma_{\mathrm{int}}=0.07$ dex. The resulting best-fit relation is therefore slightly steeper than that derived from SDSS stacks, reflecting the tendency of our high-redshift galaxies to lie along the upper envelope of the local distribution. Nevertheless, the coefficients remains consistent within the uncertainties with the local calibration of Pérez-Montero & Contini (2009).
A similar conclusion holds for the N2S2–N/O relation (right panel of Figure 3), despite its slightly larger intrinsic scatter of 0.10 dex compared to N2O2: the high-redshift galaxies – both from the MARTA sample and from the literature – are consistent with the distribution of local galaxies and the calibrations derived in the nearby Universe. The newly derived empirical fit are in good agreement with local trends, supporting the use of N2S2 as a secondary tracer of N/O in distant galaxies. The best-fit N2S2–N/O relation is
$\log(\mathrm{N/O})=(0.80\pm 0.16)\times\mathrm{N2S2}-(0.89\pm 0.08).$ (4)
@@
Figure 4: Relation between N/O and O/H for our full sample, as described in Section 2. Gold hexagons indicate MARTA galaxies with direct measurements, and light-blue triangles show the JWST auroral-line dataset. High-redshift strong-line galaxies from the literature are plotted as teal triangles, and local galaxies (both direct and strong-line measurements) as grey circles. White squares represent local galaxies binned in metallicity, showing the median N/O value in each bin. Two large blue stars indicate the median N/O values for high-redshift galaxies, with the first corresponding to $12+\log(\mathrm{O/H})<8.1$ and the second to $12+\log(\mathrm{O/H})\geq 8.1$. The black line shows the best-fit relation derived from the binned values of local galaxies using the functional form from Hayden-Pawson et al. (2022), with the best-fit parameters obtained in this work and reported in Table 1. The solar abundance values from Asplund et al. (2021) are shown with the orange Sun symbol. Finally, the pink diamonds represent higher-redshift, UV-based abundances drawn from the recent compilation by Ji et al. (2026). These sources are shown to illustrate the respective loci occupied by optical-based and UV-based abundances in this parameter space.
@@ 4.2 N/O-O/H relationship
Table 1: Best-fit parameters of the N/O–O/H relation for the local galaxy sample. [^2] The coefficients refer to equation 5. $12+\log(\mathrm{O/H})_{0}=8.43$ corresponds to $(\mathrm{O/H})_{0}=2.7\times 10^{-4}$.
@@
$1.385\pm 0.017$ $8.43\pm 0.08$ $1.8\pm 0.1$ $5.6\pm 0.9$
Figure 4 demonstrates a systematic nitrogen enhancement of the high-$z$ galaxy population relative to the local relation, particularly at lower metallicities. This is highlighted by the median N/O points as measured in the two different O/H regimes, since both lie above the local fit.
To quantitatively assess the difference between the two distributions, we compared the residuals of the high-$z$ and local samples relative to our best-fit $z=0$ relation (see Appendix C). The two distributions are statistically inconsistent with being drawn from the same parent population: the Anderson-Darling test [^3] The Anderson–Darling test is a non parametric test that, similarly to the Kolmogorov–Smirnov, compares the cumulative distributions of two samples. Unlike K-S, it assigns more weight to differences in the tails of the distributions and provides a more uniform sensitivity across quantiles. This makes it particularly suitable when the samples have very different sizes or when discrepancies in the extreme values are scientifically relevant. (AD-test, Scholz & Stephens, 1987) rejects the null hypothesis at high significance ($p<10^{-3}$). Additional tests and a full discussion are presented in Appendix C.
@@
We mostly confirm conclusions from previous works regarding nitrogen enrichment. Systems independently reported in the literature as nitrogen enhanced (e.g. the Sunburst arc, the Zhang et al. (2026) source, EXCELS objects) are also found to lie above the local trend in our analysis, and in some cases even lie outside the dispersion of our high-redshift trend. Galaxies reported to follow the local pattern (CECILIA D40, SGAS 1732) fall within the dispersion of the local population.
Stiavelli et al. (2025) classify only three of the CEERS sources as N-enhanced, noting that the remaining two are consistent with the dispersion of the local relation. With our larger statistics we find all CEERS sources to be N-enhanced with respect to the local sample, and two of them to lie above the dispersion of the high-redshift one. Our findings therefore provide context for previous discussions of individual sources in the literature, confirming the existence of both N-enhanced and non-enhanced sources, and showing how their frequency changes with metallicity.
Figure 6: Same as Figure 4, but with the SDSS density contours as the sole local reference for simplicity. The high-$z$ sample is represented with circles, with no distinction between different datasets. Left panel: the high-$z$ sample is colour-coded by $\Delta\text{MS}$, i.e. the deviation from the parametrised SFMS from Popesso et al. (2023), computed at each specific redshift. Right panel: the high-$z$ sample is colour-coded by stellar mass. The gradient angle of the $M_{*}$ trend within the plane is shown as a red arrow, visually representing the finding that N-enhanced galaxies are preferentially associated with lower stellar masses.